# Charles Gaydon

### v3

@@ IV-A Spatial and semantic diversity
Half of the dataset is found by targeted sampling of scene descriptors (47,253 patches, or 47.3%), the rest by completion sampling (52,747 patches, or 52.7%). The target minimum concentration is achieved for all descriptors except one: scenes with permanent structure points and with an antenna in the BD TOPO could not be concentrated to 1% in the train (target of $n=1000$ patches) and test (target of $n=100$ patches). Only 830 and 65 patches matching this description were found, respectively.
Figure 3: Proportional gain of each scene type in FRACTAL after sampling.
Figure 3 illustrates the proportional gain of each descriptor. The gain is most significant for rarer classes, with up to a 22-fold concentration of bridges. There is a reduction in the proportion of scenes with a slope in the $[35\leavevmode\nobreak\ m;45\leavevmode\nobreak\ m[$ range: these are common scenes, concentrated in mountainous areas, which may therefore be undersampled by the full sampling.
Figure 6 shows the distribution of FRACTAL’s patches in a small area of 4 × 6 Lidar HD tiles. The 182 patches are highlighted in yellow (targeted sampling) and cyan (completion sampling). They are scattered and cover only 1.9% of the 24 km² (52-fold reduction). Validation patches are marked with a black cross and are homogeneously distributed across both Lidar HD tiles and scene types. Two Lidar HD tiles are framed in red, which we look at in more detail in Figure 6. We observe that the targeted sampling (in yellow) over-concentrates specific landscapes such as water surfaces, specific classes such as buildings, and specific human-made structures such as parking lots, industrial buildings, and electricity pylons. 5 patches were selected by completion sampling (in cyan). Among them, 4 do not contain any scene type flag. This illustrates the scarcity of more complex scenes in geographical data and the need to explicitly target them to obtain a challenging and diverse dataset.
IV-B File Structure
Figure 4: Structure of files and directories in FRACTAL.
Figure 4 illustrates the dataset’s file structure. The point clouds are organized by set (train, val, and test), in 100 subdirectories of 1000 point clouds each. The naming convention for data patches is shown in the figure, where tile_id gives the X and Y kilometer northwest coordinates of the Lidar HD tile in the Lambert 93 projection (EPSG:2154), and patch_id is a unique, arbitrary patch identifier. For instance, patch TEST-0744_6246-006804306.laz belongs to the test set, was extracted from Lidar HD tile whose top left coordinates are $X=0744$ and $Y=6246$, and its unique identifier is $006804306$.
The geometries of all patches and their descriptions are provided in a metadata file: fractal.gpkg.
IV-C Limitations and recommendations
FRACTAL is sampled from an existing open ALS archive, and shares its limitations. The dataset only includes data that underwent human corrections and auditing (i.e. “Optimized Lidar HD”), however, the specification of the Lidar HD classification includes a tolerance for errors. This tolerance is class specific: for instance, a perfect 100% recall for individual buildings is not mandatory as explained in Section II-A). Beyond actual errors, the very definition of “unclassified” points leads to potential confusions with other classes which we discuss in Section V-A. Considering the size of the dataset, these imperfections are expected to have a negligible impact for model evaluation, but users should be aware that they exist when inspecting the data.
Point clouds are colorized with asynchronous aerial imagery, which can lead to color artifacts (typically for moving objects such as vehicles) and misalignment between colors and shapes that can blur segmentation boundaries (e.g., roof points colored by a nearby tree). Mitigation strategies such as color augmentation and randomly dropping color dimensions may help reduce models’ reliance on color information.
Using FRACTAL to train models for production purposes should be done with caution. Indeed, the dataset covers 5 spatial domains from 5 southern regions of metropolitan France. While large and diverse, it only covers a fraction of the French territory and is not representative of its full diversity regarding either landscapes or human-made structures. Furthermore, domain shifts are frequent for aerial images due to different acquisition conditions and downstream data processing. When using models trained on the dataset to predict a classification in unseen regions, one should make the adequate verification and not assume its capacity to generalize. Similarly, considering new Lidar sensors should be done carefully. ALS data of comparable point densities (about 40 pts/m²) are expected to have consistent geometric characteristics, but users should still assess the model’s accuracy when predicting from alternative 3D data sources.
Figure 5: Patches in FRACTAL on a subset area of 4 × 6 Lidar HD tiles. Patches selected via targeted sampling (yellow) are more present in complex urban areas. Patches from completion (cyan) sampling are homogeneously distributed.
@@ IV-C Limitations and recommendations
Figure 6: Enlarged view of 2 Lidar HD tiles showing the difference between targeted sampling and completion sampling. 25 patches are selected from the 800 possible patches, most of which contain scenes of particular interest.
V Baseline evaluation
V-A Task and metrics
@@
FRACTAL is a benchmark dataset for semantic segmentation of ALS scenes into 7 semantic classes: other, ground, vegetation, building, water, bridge, and permanent structure. This reduced nomenclature is adapted from the Lidar HD nomenclature, which has 11 classes (described in Appendix LABEL:appendix:lhd_nomenclature). Low, medium, and high vegetation are grouped together since they only differ in height above ground, artefact and synthetic points are simply filtered out, and the “Unclassified” ASPRS class is renamed "other" for clarity. This correspondence between the two nomenclatures is summarized in Figure IV.
TABLE IV: Relation between the Lidar HD classification nomenclature and its adaptation in FRACTAL.
Lidar HD $\rightarrow$ FRACTAL
Unclassified $\rightarrow$ Other
Ground $\rightarrow$ Ground
Low vegetation Vegetation
Medium vegetation $\rightarrow$
High vegetation
Buildings $\rightarrow$ Building
Water $\rightarrow$ Water
Bridge deck $\rightarrow$ Bridge
Permanent structures $\rightarrow$ Permanent structure
Artefact -
Synthetic
We adopt the evaluation practices of similar Lidar datasets (e.g. [12]) and use the mean IoU (mIoU) as our main metric. Importantly, we decide not to exclude class other from metric computations. This class is semantically ill-defined and may include points with blurred boundaries with adjacent classes such as building, ground, and vegetation. However, it also includes well-defined objects such as vehicles and containers. Such objects present a known challenge to segmentation methods, and we believe it is important to account for them in model evaluation.
@@
For consistency with other benchmarks, we also report Overall Accuracy (OA), and precision, recall, and F1 score (for each class and macro-averaged). All metrics should be computed from the accumulated point confusion matrix over all point clouds from the test set, which we also report as part of our baseline evaluation.
V-B Baseline model
@@
The baseline performance is established using Myria3D [24], a 3D deep learning library developed at IGN. Myria3D leverages the Pytorch-Lightning framework [25], and Pytorch-Geometric [26], a powerful library for 3D deep learning.
V-B1 Neural architecture
@@
Myria3D was explicitly built for the semantic segmentation of Lidar HD data, and scalability informed its design, including the choice of its neural architecture: RandLa-Net [27]. Since the PointNet++ architecture [10] succeeded to the ground-breaking PointNet [9] to operate directly on unordered point clouds, there were many attempts to improve over point-based architectures, characterized by PointNet-like operations hierarchically organized in a U-shaped architecture. Conceptually simple, RandLa-Net makes some interesting additions to PointNet++. It uses a lightweight module for local spatial encoding and achieves performance gains thanks to random sampling. Importantly, authors demonstrated its performance in large-scale outdoor Lidar benchmark datasets like SemanticKITTI [28] and Semantic 3D [29].
V-B2 Data processing and training hyperparameters
In our experiments, we kept most defaults setting from Myria3D, which were described in extenso in a previous work. We thus refer the reader to Section Benchmark Models in [11] for details on the processing of colorized Lidar HD tiles as well as training hyperparameters (optimizer, learning rate, scheduler, early stopping, etc.).
V-B3 Infrastructure
Training is conducted with 6 NVIDIA Tesla V100 GPUs, each equipped with 32 GB of memory, using Pytorch-Lightning’s distributed data parallel (ddp) strategy. We log all metrics using Comet, a machine learning experiment tracking tool [30]. With this configuration the approximate learning time is 30 minutes per epoch.
V-C Experimental results
We report the results obtained using the 80,000 patches for training and 10,000 patches for validation, and testing on the remaining 10,000 patches of the dataset.
V-C1 Quantitative results
@@
The baseline achieves a test mIoU of 77.5% and a test OA of 96.1%. Figure 7 details the IoU results per class. We observe highly accurate results for the three most common classes: ground, vegetation, and building have an IoU above 90%, with a maximum of 93.8% for vegetation. Similarly, class water has an IoU of 90.1%, despite its extreme initial rarity (0.6%) in the sampling area. For classes bridge and permanent structure, also initially rare (0.01% each), the model achieves decent performance with an IoU above 60%. Compared to the six other classes, the baseline underperforms for class other, but still achieves an IoU of 47.5%, which is remarkable considering the fuzzy definition of this class. Based on our own experience with Lidar HD semantic segmentation, these performances can not be achieved without the class rebalancing that happens in the sampling of FRACTAL. For an in-depth performance assessment, all metrics (IoU, precision, recall and F1 score) are reported in Appendix LABEL:appendix:metrics.
othergroundvegetationbuildingwaterbridgeperm. struct.005050505010010010010047.547.547.547.591.991.991.991.993.893.893.893.890.490.490.490.490.190.190.190.165.265.265.265.263.563.563.563.5IoU (%)
Figure 7: Baseline per-class test IoUs.
Figure 8 shows the baseline confusion matrix on the test set, normalized by both rows (a) and columns (b). Vegetation, ground, and building have low relative confusion, as expected from their high IoU. Permanent structure has its highest confusion with vegetation, due to the vertical nature and intricate geometries shared by both classes e.g., between pylons and trees. High interclass confusion is also found between classes that may have imprecise geometric boundaries. For example, bridge points are misclassified as ground, of which they are often an extension. Points of class other have high confusion with vegetation and ground points. This is mainly due to the Lidar HD classification specification, which favors specificity over recall for the ground and vegetation classes. As a result, the model may be penalized for accurately identifying ground and vegetation whose target class is other. Note that with only 0.6% of all points in FRACTAL, class other is a small minority, and these errors have a negligible effect on the measured performance for the ground and vegetation classes.
Figure 8: Baseline test confusion matrices normalized by rows (a) and columns (b). Diagonal values in row-normalized matrix correspond to recall. Diagonal values in column-normalized matrix correspond to precision.
V-C2 Qualitative results
Finally, Figure 9 presents sample predictions of the baseline model for five types of Lidar scenes: a) OTHER_PARKING, b) WATER and BRIDGE, c) URBAN, d) BUILD_GREENHOUSE, and e) HIGHSLOPE1. The 5 patches are selected at random based solely on their scene descriptors, among test patches with at least 10k points (i.e. at least 4 pts/m²). We present the inference results without cherry-picking.
Figure 9: Input point cloud, target classification, and predictions of the baseline model for a subset of test patches. Patches are selected at random and without cherry-picking, to match the following scene types: a) OTHER_PARKING, b) WATER and BRIDGE, c) URBAN, d) BUILD_GREENHOUSE, and e) HIGHSLOPE1. Color scheme is: other (white), ground (orange), vegetation (green), building (red), water (cyan), bridge (yellow), permanent structure (purple). Predictions errors are circled in red.
VI Conclusion
@@
We present FRACTAL, an ultra-large-scale benchmark dataset for the 3D semantic segmentation of ALS point clouds. The dataset is based on high quality open data from the French Lidar HD nationwide acquisition program. It is a distillation of a larger initial area of 17,440 km² in five French regions. We show that a simple targeted sampling with spatial stratification at all levels effectively preserves the spatial and semantic diversity of regional-scale volumes of Lidar data. FRACTAL’s size is compatible with deep learning research practices, and it is the largest Lidar benchmark dataset to date, with 9261 million points in 100,000 point clouds and a total span of 250 km². While other ALS benchmark datasets typically cover a single urban area, the large diversity of urban and rural landscapes in FRACTAL is representative of the challenges of 3D semantic segmentation for land monitoring. Thanks to class rebalancing, our methodology opens the possibility of robust assessment of semantic segmentation performance, even for initially rare classes water, bridge, and permanent structure. The baseline evaluation of a 3D neural network (RandLa-Net) further demonstrates the quality of the dataset for model benchmarking. We invite the research community to benchmark both state-of-the-art and novel methods against this dataset. We hope that FRACTAL will advance the field of deep learning for airborne Lidar and ultimately benefit public Lidar-based 3D mapping programs for land monitoring.
@@ Acknowledgements
The authors thank Léa Vauchier for her code reviews of the data engineering code, and Marouane Zellou for maintaining the in-house high-performance computing server, and Matthieu Porte and Anatol Garioud for reviewing this data paper. The authors also thank Anatol Garioud for paving the way for the open release of deep learning datasets for land monitoring at IGN with the FLAIR dataset [31].
@@
Appendices
A Specifications of the Lidar HD classification.
RGB Value Name Content
0,255,0 3 Low vegetation Trees, shrubs and low vegetation (e.g. bushes, ferns, reeds, etc.). Vegetation at ground level (less than 20 cm high, typically grass) is classified as vegetation only if there are enough ground points locally for ground modeling. The class also includes cultivated trees (orchards, vineyards, field crops, etc.). Vegetation is further divided into 3 classes based on height above ground: low (< 0.5m), medium (between 0.5 m and 1.5 m), and high ($\geq{1.5\leavevmode\nobreak\ m}$) vegetation.
0,193,0 4 Medium vegetation
0,134,0 5 High vegetation
255,255,0 17 Bridge deck A bridge is an engineering structure passing over one or more elements of the road, rail or waterway network. Point on bridge decks are included, while structural elements such as piers and parapets are assigned to the "Unclassified" class. Very high structural elements ($\geq$ 5 m above deck level) such as piers and abutments are classified as “Permanent structures”. Tunneled passages (including nozzles, which are openings in the ground generally to allow water to drain) are considered part of the ground and therefore excluded.
64,0,128 65 Artefact All points that do not correspond to an actual object or terrain.
Refer to the Lidar HD product description [19] for more specifications.
B Sensors used to acquire Lidar HD data in each of the 8 50 $\times$ 50 km blocks represented in FRACTAL.
Subcontractor(s)
Block Sensor Acquisition Classification
GN CityMapper 2H: Hyperion 2+ APEI + Avineon Avineon
MQ RIEGL VQ1560 II-S Eurosense + SFS Eurosense + SFS
MP RIEGL VQ-1560 II Eurosense Sintégra
PK RIEGL VQ1560 II Eurosense Eurosense
PO Leica CityMapper-2 Avineon + APEI Avineon + APEI
PP Leica CityMapper-2 Avineon + APEI Eurosense + SFS
QO RIEGL VQ780 II-S Sintégra + Bluesky Sintégra + Bluesky
UT RIEGL VQ780 II-S Sintégra + Bluesky Avineon + APEI
N.B. Data acquisition and its classification were often conducted by distinct subcontractors, which we also report.
C Number of points by class in the sampling area, in the dataset, and in the train, val and test sets.

### v4

@@ 4.3 Spatial and landscape diversity in FRACTAL
Figure 3: Enlarged view of two Lidar HD tiles and their sampled patches, most of which contain scenes of particular interest.
The target proportions for each scene type are specified in Tab. 2. They are chosen so that about 20% of sampled patches consist of specific landscapes, such as seashores and forests, while about 30% of them contain specific structures and objects, such as highways and water surfaces. Half of the dataset is found by targeted sampling (47,253 patches, or 47.3%), then completed (52,747 patches, or 52.7%) until the target dataset size is reached. The target minimum proportion is achieved for all scene types except one: scenes with permanent structure points and with an antenna in the BD TOPO could not amount to 1% in the train set (target of $n=1000$ patches, 830 were found) and test set (target of $n=100$ patches, 65 were found).
Fig. 3 presents a visualization of the sampling in two Lidar HD tiles, where 25 patches are selected from the 800 possible patches. Targeted sampling (yellow) over-concentrates specific landscapes such as water surfaces, specific classes such as buildings, and specific human-made structures such as parking lots and electricity pylons. On the other hand, four of the five patches selected by completion sampling (cyan) do not contain any predefined scene type. This illustrates the scarcity of complex scenes in geographical data and the need to explicitly target them to obtain a challenging and diverse dataset. See Sec. B.3 for a similar illustration at a larger scale.
Fig. 4 illustrates how our sampling increases the presence of targeted scene types in FRACTAL. The gain is most important for rarer classes, with up to a 22-fold concentration of scenes with a bridge. All scene types of interest end up with better representation in FRACTAL; except for high-slope scenes: most of them are spatially concentrated in mountainous areas, and thus slightly undersampled by completion sampling.
Figure 4: Proportional change of proportion of scene types in FRACTAL after sampling.
Additional figures are reported in the appendices: proportions of each semantic class (B.1) and prevalence of scene types in train, val, and test sets (B.2).
4.4 Data extraction and colorization
Point clouds are made available in the LAZ 1.4 format defined by the American Society for Photogrammetry and Remote Sensing (ASPRS) [23]. The following dimensions are relevant for semantic segmentation: xyz coordinates in Lambert 93 spatial reference system (EPSG:2154), intensity (16-bits encoding), return number and number of returns, scan angle, and near-infrared, red, green and blue (16-bits encoding). The Lidar HD classification nomenclature, detailed in Appendix D.1, is left intact in the data patches.
4.5 Limitations and recommendations
FRACTAL is sampled from an open ALS archive, and shares its limitations. The dataset only includes data that underwent human corrections and auditing (i.e. “Optimized Lidar HD”), however, the specification of the Lidar HD classification tolerates some errors. This tolerance is class specific; for instance, a perfect 100% recall for individual buildings is not mandatory (as previously stated, see Section 3.1).
Beyond errors, the very definition of “unclassified” points leads to potential confusions with other classes. This class is semantically ill-defined and may include points with blurred boundaries with adjacent classes such as building, ground, and vegetation. These imperfections are expected to have a negligible impact for model evaluation, but users should be aware that they exist when inspecting the data.
Point clouds are colorized with non-simultaneous aerial imagery (as explained in Section 3.2), which can lead to color artifacts (typically for moving objects such as vehicles) and to misalignment between colors and shapes, which can blur object boundaries (e.g., roof points colored by a nearby tree). We observed only marginal impact on segmentation performances; overall, the colorization of the point clouds benefits segmentation models.
Using FRACTAL to train models for production should be done with caution. The dataset covers 5 spatial domains from 5 southern regions of metropolitan France. While large and diverse, it reprents only a fraction of the French territory and is not representative of its full diversity regarding either landscapes or human-made structures. Furthermore, domain shifts are frequent in aerial images due to different acquisition conditions and downstream data processing. When using models trained on the dataset to predict a classification in unseen regions, one should make the adequate verification and not assume its capacity to generalize. Similarly, considering new Lidar sensors should be done carefully. ALS data of comparable point densities (about 40 pts/m²) are expected to have consistent geometric characteristics, but users should still assess the model’s accuracy when predicting from alternative 3D data sources.
5 Experimental results
5.1 Task and metrics
@@
FRACTAL is a benchmark dataset for semantic segmentation of ALS scenes into 7 semantic classes: other, ground, vegetation, building, water, bridge, and permanent structure. This reduced nomenclature is adapted from the Lidar HD nomenclature, which has 11 classes (see Appendix D.1). Low, medium, and high vegetation are grouped together since they differ only in height above ground. Artifact and synthetic points are simply filtered out. The “unclassified” ASPRS class is renamed ”other” for clarity. See Sec. D.2 for a table of correspondence between the two nomenclatures.
We adopt the evaluation practices of similar Lidar datasets (e.g., DALES [24]) and use the mean Intersection over Union (mIoU) as our main metric. The ”other” class is imperfect (see Section 4.5), but we decide not to exclude it from metric calculations since it contains well-defined objects that are relevant to scene segmentation (e.g., boats, containers). Considering the size of the dataset, imperfections of the ”other” class are expected to have a negligible impact in model evaluation.
@@
For consistency with other benchmarks, we also report Overall Accuracy (OA). All metrics should be computed from the confusion matrix accumulated over all point clouds from the test set.
5.2 Baseline model
@@
A baseline performance is established using Myria3D [7]. Myria3D is a 3D deep learning library developed at IGN, which leverages Pytorch-Lightning [4], and the the 3D deep learning library PyTorch-Geometric [5]. It was explicitly built for the semantic segmentation of Lidar HD data. Scalability informed its design, including the choice of its neural architecture: RandLa-Net [9].
@@
Since the PointNet++ architecture [21] succeeded to the ground-breaking PointNet [20] to operate directly on unordered point clouds, there were many attempts to improve over point-based architectures, characterized by PointNet-like operations hierarchically organized in a U-shaped architecture. Conceptually simple, RandLa-Net makes some interesting additions to PointNet++. It uses a lightweight module for local spatial encoding and achieves performance gains thanks to random sampling. Importantly, its authors demonstrated its segmentation accuracy on large-scale outdoor Lidar benchmark datasets like SemanticKITTI [1] and Semantic 3D [8].
To produce the first experimental results on the benchmark, we kept the default hyperparameters of [3d-deep-learning-library]. Details about the processing of colorized Lidar HD tiles, about hyperparameters (optimizer, learning rate, scheduler, early stopping, etc.), and about infrastructure, are given in Sec. C.4.
5.3 Results
Table 4: Baseline test IoUs and OA
Class IoU OA
other 47.5 54.9
ground 91.9 97.7
vegetation 93.8 95.6
building 90.4 93.7
water 90.1 92.6
bridge 65.2 96.1
permanent structure 63.5 76.6
Macro Average 77.5 86.7
@@ 5.3 Results
The baseline model achieves a test mIoU of 77.5% and a test OA of 96.1%. Tab. 4 reports the IoU and OA for each class. We observe highly accurate results for the three most common classes: ground, vegetation, and building have an IoU above 90%, with a maximum of 93.8% for vegetation. The ”water” class has an IoU of 90.1%, despite its extreme initial rarity (0.6%) in the area of interest. For the ”bridge” and ”permanent structure” classes, also initially rare (0.01% each), the model achieves decent performance with an IoU above 60%. The baseline underperforms for the ”other” class compared to the six other ones, but still achieves an IoU of 47.5%, which is remarkable considering the fuzzy definition of this class.
Additional metrics (precision, recall, and F1 score) are reported in Appendix C.1. Confusions matrices are also given (Appendix C.2). Finally, the qualitative performance of the baseline model can be assessed from examples of model predictions (Appendix C.3).
6 Conclusion
@@
We present FRACTAL, an ultra-large-scale benchmark dataset for the 3D semantic segmentation of ALS point clouds. The dataset is based on high quality open data from the French Lidar HD acquisition program. It is a distillation of a larger initial area of 17,440 km² in five French regions. We show that a simple targeted sampling with spatial stratification at all levels preserves the diversity of regional-scale volumes of Lidar data. FRACTAL’s size is compatible with deep learning research practices, and it is the largest Lidar benchmark dataset to date, with 9,261 million points in 100,000 point clouds and a total span of 250 km². While other ALS benchmark datasets typically cover a single urban area, the large diversity of urban and rural landscapes in FRACTAL is representative of the challenges of 3D semantic segmentation for land monitoring. Thanks to class rebalancing, our methodology opens the possibility of robust assessment of semantic segmentation performance, even for the initially rare classes water, bridge, and permanent structure. The baseline evaluation of a 3D neural network (RandLa-Net) further demonstrates the quality of the dataset for model evaluation. We invite the research community to benchmark both state-of-the-art and novel methods against this dataset. We hope that FRACTAL will advance the field of deep learning for airborne Lidar and ultimately benefit public Lidar-based 3D mapping programs for land monitoring.
@@ Acknowledgements
The authors thank Léa Vauchier for her code reviews of the data engineering code, and Marouane Zellou for maintaining the in-house high-performance computing server, and Matthieu Porte and Anatol Garioud and Nicolas Audebert for reviewing this data paper. The authors also thank Anatol Garioud for paving the way for the open release of deep learning datasets for land monitoring at IGN with the FLAIR dataset [6].
@@
Appendices of the FRACTAL datapaper
Appendix A Data representation for patch sampling
A.1 Considerations for data cataloguing
We present the data representation used for sampling, which takes the form of a PostGIS database. Cataloging data boils down to two steps: listing and describing, which we illustrate in Sec. A.2. We start by regularly dividing the area of interest into $50\times 50$ m patches. We then describe all patches with two sets of descriptors: Lidar descriptors and vector descriptors. Lidar descriptors are derived from the Lidar data patches themselves: number of points in each semantic class, elevation (above sea level), and elevation gain. Vector descriptors are derived by cross-referencing the patches with the BD TOPO® [13], a 3D vector description of the French territory and its infrastructures (buildings, roads, etc.). We include flags indicating the presence of the following objects: greenhouses, antennas, pylons, highways, parking lots, water surfaces, greenhouses, pylons, antennas, and industrial, commercial, or agricultural buildings.
Cataloguing with a PostGIS database gives us the full power of SQL syntax to select data and to further characterize each Lidar patch with bespoke SQL queries that combine simple descriptors into more complex ones (see Tab. 2).
A.2 Data cataloguing
The figure shows how a single Lidar HD Tile of 1 km x 1 km is divided into many smaller patches which are then described with vectore features and point cloud features.
Figure 5: Data cataloguing boils down to a) listing and b) describing all Lidar patches, prior to sampling.
Appendix B Dataset content
B.1 Number of points by class in the area of interest, in the dataset, and in the train, val and test sets